Hello I made a mistake earlier and was asked to repost this. I just can not deal with word problems. I always mess them up please help me solve this.
Everton college $1625 for an order of 45 calculators. the store paid $10 for each scientific calculator, The others all graphing calculators cost the store $57 each. How many of each type were ordered?
Thank you again...

Answer by stanbon(48558)

Everton college paid $1625 for an order of 45 calculators. The store paid $10 for each scientific calculator, The others all graphing calculatos cost the store $57 each. How many of each type were ordered?
-------------------------
Let # of scientific calculators be "x" ; value of these is 10x dollars
--------------
# of graphing calculators is "45-x" ; value is 57(45-x)= 2565 -57x dollars
--------------
EQUATION:
value + value = 1625 dollars
10x + 2565-57x = 1625
-47x = -1240
x = 26.38..
---------
Comment: Something is wrong with your data. Since the value
of "x" is the number of scientific calculators, the value
of x should be a counting number, not a decimal.
Please check your data.
--------------
Cheers,
Stan H.
shou